HIIII i`m Varun, can anybody tell me how to do this one : "Find the area enclosed by y=g(x), x=1, and x=15 where g is inverse of f given by f(x) = x^3+3x+1"????    Answer is 18....

Grade:12

1 Answers

Pratik Tibrewal
10 years ago
follow this answer from the last post...

x^3+3x+1=1
x=0

x^3+3x+1=15
(x-2)(x^2+2x+7)=0
x=2
=>f^(-1) (1)=0 & f^(-1) (15)=2
=>15*2-[2^4/4+3*2^2/2+1*2]
=18


Thanks and Regards,
Pratik Tibrewal,
askiitians faculty,
BTech